OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x − 2y = 16?

OpenStudy (anonymous):

wherre are order pair..?

OpenStudy (anonymous):

(-5,-3) (-5,3) (6,-4) (6,4)

OpenStudy (anonymous):

lets find out... 4x - 2y = 16 (-5,-3) 4(-5) - 2(-3) = 16 -20 + 6 = 16 - 14 = 16 (incorrect) Not this one (-5,3) 4(-5) - 2(3) = 16 -20 - 6 = 16 - 26 = 16 (incorrect) nope...not this one either (6,-4) 4(6) - 2(-4) = 16 24 + 8 = 16 32 = 16 (incorrect) nope (6,4) 4(6) - 2(4) = 16 24 - 8 = 16 16 = 16 (correct) Here it is...it would have to be the last one I tried...lol

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

its (6,4)